In an economy C= 200 + 0.5 Y is the consumption function where C is the 

consumption expenditure and Y is the national income. Investment 

expenditure is ₹ 400 crores.

Is the economy in equilibrium at an income level ₹1500 crores? Justify your 

answer.


1
Expert's answer
2021-02-26T10:36:18-0500

"No, the \ economy\ is\ not \ at\ a\ state \ of \ equilibrium\ at\ an\ income\ level \\of \ 1500\ crores\ given\ consumption\ function,"


"C = 200 + 0.5Y"

"Investment\ expenditure(I)=400\ crores"

"At\ the\ equilibrium\ level :"

"Y=C+I"


"Substituting\ the\ values\ from \ the \ question:"


"Y=({200+0.5Y})+400\\
Y-0.5Y=600\\
0.5Y=\frac{600}{0.5}=1200\\"


"The \ equilibrium \ level \ of\ income\ is\ 1200\ crores.\\
So\ the\ economy\ will\ be\ equilibrium\ only \ if\ Y=1200.\\
The \ given \ income \ is \ 1500\ crores\ which\ is\ greater\ than \ the\ equilibrium\ level\\
 of\ income\, therefore\ the \ economy\ is\ not\ at\ equilibrium."